# Multivariate Temporal Regression at Scale: A Three-Pillar Framework Combining ML, XAI, and NLP

### v1

@@ VII Conclusion
This study introduces a novel framework for enhancing data quality in high-dimensional temporal regression tasks by integrating machine learning (ML) and explainable AI (XAI). By addressing the inherent challenges of temporal data—such as evolving variable relationships, noise in time-series inputs, and domain-specific contextual dependencies—the proposed methodology bridges the gap between automated data refinement and human-centric interpretability.
Central to this approach is the transformation of multivariate temporal data into structured 2D arrays, enabling image-based architectures like ResNet and ResNext to detect patterns obscured by noise or redundancy. The integration of XAI tools (SHAP, LIME) ensures that pruning decisions are both data-driven and aligned with domain expertise. For instance, global heatmap aggregations revealed time-specific biases in sensor readings, enabling targeted calibration of critical features.
The framework demonstrates significant improvements in computational efficiency (25% reduction in training time) and predictive accuracy (25% MAE improvement) compared to baseline methods, even across heterogeneous hardware platforms. By preserving critical temporal dynamics while eliminating irrelevant features, the methodology balances scalability with contextual relevance—a critical requirement for applications like agricultural yield forecasting, where seasonal variability and sensor inconsistencies dominate.
Future work will extend this framework to real-time temporal systems and explore its applicability to other domains with dense time-dependent data, such as climate modeling and industrial IoT. This research underscores the transformative potential of data-centric AI pipelines that prioritize temporal coherence and human collaboration to unlock robust, scalable solutions.

### v2

@@ VII Conclusion
This paper presents a three-pillar framework integrating machine learning (ML), explainable AI (XAI), and future NLP-driven contextual validation to address data quality challenges in high-dimensional temporal regression. The framework automates noise reduction and bias correction while preserving critical feature dynamics by repurposing image-based architectures (e.g., ResNet) to analyze structured 2D temporal arrays. XAI tools (SHAP, LIME) ensure interpretable pruning decisions, linking data refinements to domain knowledge. Evaluated on agricultural and synthetic datasets, the approach achieves 25% reductions in training time and 25% improvements in predictive accuracy (e.g., MAE), with hardware-agnostic scalability across diverse platforms. This work establishes a foundational pipeline for accelerating data-driven discovery in temporal regression tasks. By reducing the time to identify critical feature relationships, the framework enables stakeholders to act sooner on insights, laying the groundwork for measurable real-world impact. Follow-up studies will focus on longitudinal validation and integration with NLP pipelines to further shorten the path from data to decision.
Acknowledgment
The authors acknowledge the use of Qwen [21], a large language model developed by Alibaba Cloud, for assisting in the research, analysis, and writing phases of this work.